A fan blade is initially rotating an angular speed of 4.6 rpm. It slows down and eventually comes to rest in a time of 32 seconds after turning through a total of 8.8 revolution. Find a.) The angular speed in rad/s. b.) The average angular velocity ang c.) The average angular acceleration.
1
Expert's answer
2020-06-09T13:20:14-0400

Initial angular velocity is "\omega_o=4.6\times \frac{2\pi}{60}=0.48\:rad/s"

Final angular velocity is "\omega_f=0\:rad/s"

Final time "t=t_f=32s" and angle traversed is "\theta=8.8\times2\pi=55.29 \:rad"


(a).As we have not provided at which time we need to find the angular speed, thus

we assume time t is arbitrary, hence

"\omega=\omega_o+(-\alpha) t\implies \omega=\omega_o-\alpha t\:rad/s"

As angular acceleration is decreasing.


(b). Since, average angular velocity is given by

"\omega_{av}=\frac{\Delta \theta}{\Delta t}"

Thus, on plugin the value we get,

"\omega_{av}=\frac{55.29-0}{32-0}=1.72 \: rad/s"

(c). Since,

"\alpha_{av}=\frac{\Delta \omega}{\Delta t}=\frac{\omega_f-\omega_o}{t_f-t_o}"

Hence, on plugin the value we get

"\alpha_{av}=\frac{0-0.48}{32-0}=-0.015 \:rad/s^2"